# Generating Reservoir State Descriptions with Random Matrices

### v2

@@ 3.4 Quantum Circuit Reservoir Computing on the Mackey-Glass Series
4 Practical Realisation
The results above demonstrate the potential of this reservoir measurement approach. However, its practical application to real hardware remains a question. Currently, performing expectations over hundreds of measurements using quantum devices is not feasible. Nonetheless, there are ways to overcome this challenge. In many real devices, scientists conduct measurements further away from the driving mechanism, allowing multiple measurements to be performed without disturbing the primary dynamics. In these cases, the driving signal is often repeated to obtain the required statistics, resulting in a non-real-time workflow. Looking to the future, just as we utilize chips limited by a bit architecture, it’s intriguing to imagine that this might extend to reservoir computing, where small devices like stacks of spin lattices are designed to perform hundreds of measurements over thousands of ensembles simultaneously.
5 Conclusion and Outlook
@@
We have introduced a new approach to measuring quantum reservoir computers using random matrices. Our method can be scaled to arbitrary state dimensions and performs well in test cases utilizing two quantum reservoirs, a Heisenberg spin chain, and a quantum circuit. While the results shown here correspond to quantum systems, any reservoir computer paradigm could be substituted as long as the measurement problem can be phrased in matrix operations. As a first investigation, we showed the spectrum and observable range of the random matrices on random system states. We found that partial traces over the entire system may improve the diversity of the observables that can be used in reservoir computing. We then demonstrated the efficacy of our method on the problem of time-series prediction for a sine wave and the Mackey-Glass time series, as well as data interpolation on historical Dow30 data. Our results suggest that random matrices work well for different reservoirs for state description. One avenue for future research is the study of much more strongly correlated spin systems or alternative reservoirs with a better learning capacity. Furthermore, larger reservoirs should be probed to better understand the influence of correlations spanning more degrees of freedom, possibly also in experiments.

### v3

@@ Introduction
Conclusion and Outlook
@@
We presented a new approach to measuring quantum reservoir computers using random matrices. Our method is scalable to arbitrary state dimensions and performs well in test cases using two quantum reservoirs, a Heisenberg spin chain, and a quantum circuit. While the results shown here correspond to quantum systems, any reservoir computing paradigm could be substituted as long as the measurement problem can be formulated in terms of matrix operations. As a first investigation, we showed the spectrum and observable range of random matrices on random system states. We found that partial traces over the entire system may improve the diversity of observables that can be used in reservoir computation. We then demonstrated the effectiveness of our method on the problem of time series prediction for a sine wave and the Mackey-Glass time series, as well as data interpolation on historical Dow30 data. Our results suggest that random matrices work well for different reservoirs for state representation. However, its practical application to real hardware remains an open question. Currently, it is not feasible to fulfill expectation values over hundreds of measurements with quantum devices. However, there are ways to overcome this challenge. In many real devices, scientists perform measurements far away from the driving mechanism so that multiple measurements can be made without perturbing the primary dynamics. In these cases, the drive signal is often repeated to obtain the required statistics, resulting in a non-real-time workflow. Looking to the future, just as we use chips that are limited by a bit architecture, it’s intriguing to imagine that this could be extended to reservoir computing, where small devices such as stacks of spin lattices are designed to perform hundreds of measurements over thousands of ensembles simultaneously. One avenue for future research is to investigate much more highly correlated spin systems or alternative reservoirs with better learning capabilities. In addition, larger reservoirs should be explored to better understand the influence of correlations spanning more degrees of freedom, possibly also in experiments. Another key element to be studied is the influence of quantum phenomena such as entanglement and the phase space used on the performance of the storage. The aim is to determine whether true quantum phenomena influence reservoir performance. While for a certain QRC architecture an influence of quantum properties on the reservoir performance has been observed (Götting, Lohof, and Gies 2023), these results have to be investigated for the architectures of this paper in future work.
Acknowledgments
C.H, S.T and T.F. acknowledge financial support from the German Funding Agency (Deutsche Forschungsgemeinschaft DFG) under Germany’s Excellence Strategy EXC 2075-390740016. C.H, S.T and T.F. acknowledge financial support from the German Funding Agency (Deutsche Forschungsgemeinschaft DFG) under the Priority Program SPP 2363, “Utilization and Development of Machine Learning for Molecular Applications - Molecular Machine Learning” Project No. 497249646. The authors acknowledge support by the state of Baden-Württemberg through bwHPC and the German Research Foundation (DFG) through grant INST 35/1597-1 FUGG. The authors acknowledge support from the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) Compute Cluster grant no. 492175459.
Data availability statement
The files to reproduce the findings of this study are openly available in the data repository of the University of Stuttgart (DaRUS) in the Dataset ”Replication Scripts for: Generating Quantum Reservoir State Representations with Random Matrices”: https://doi.org/10.18419/DARUS-4742.